# PV-S3: Advancing Automatic Photovoltaic Defect Detection using Semi-Supervised Semantic Segmentation of Electroluminescence Images

### v1

@@ Abstract
Photovoltaic (PV) systems allow us to tap into all abundant solar energy, however they require regular maintenance for high efficiency and to prevent degradation. Traditional manual health check, using Electroluminescence (EL) imaging, is expensive and logistically challenging making automated defect detection essential. Current automation approaches require extensive manual expert labeling, which is time-consuming, expensive, and prone to errors. We propose PV-S3 (Photovoltaic-Semi Supervised Segmentation), a Semi-Supervised Learning approach for semantic segmentation of defects in EL images that reduces reliance on extensive labeling. PV-S3 is a Deep learning model trained using a few labeled images along with numerous unlabeled images. We introduce a novel Semi Cross-Entropy loss function to train PV-S3 which addresses the challenges specific to automated PV defect detection, such as diverse defect types and class imbalance. We evaluate PV-S3 on multiple datasets and demonstrate its effectiveness and adaptability. With merely 20% labeled samples, we achieve an absolute improvement of 9.7% in IoU, 29.9% in Precision, 12.75% in Recall, and 20.42% in F1-Score over prior state-of-the-art supervised method (which uses 100% labeled samples) on UCF-EL dataset (largest dataset available for semantic segmentation of EL images) showing improvement in performance while reducing the annotation costs by 80%.
@@ 1 Introduction
The increasing global demand for clean and sustainable energy sources has highlighted the importance of solar energy, particularly photovoltaic (PV) systems, in the renewable energy landscape [23, 28, 43]. While PV modules are essential for converting sunlight into electricity, their efficiency and reliability are often compromised by various defects, ranging from manufacturing imperfections to environmental stressors [22, 33, 19, 32, 15, 52, 20, 26, 6, 16, 53, 27]. These defects reduce the overall efficiency and the life cycle while impacting the financial viability of PV systems. Various imaging techniques such as Infrared Thermography (IRT), Photoluminescence (PL) Imaging, and Electroluminescence (EL) imaging are currently used, however, EL imaging remains a key technique due to its high resolution and detailed electrical property information [48, 21, 30, 5, 11, 50, 31]. Traditionally, defect detection in PV modules has relied on manual inspection of EL images [7, 18, 17]. Manual inspection however requires certain expertise while being labor-intensive and prone to human error. Recent expansion of PV installations increased the number of images that need to be analyzed thus exacerbating the problem. In long run automation is the key to address all these issues.
@@
Application of artificial intelligence (AI to bring automation has been widely accepted in the energy industry [38, 10, 45, 25] and owing to above challenges, application for automated defect detection in PV modules has also gained attention [1, 2, 3, 21, 44]. AI-based approaches, including binary classification for defective/non-defective, multi-class classification for identifying types of defects, and object detection for locating defects, pave the way for comprehensive analysis of PV systems [24, 54]. Among these, semantic segmentation stands out for its ability to delineate and classify each pixel of an image according to defect type, offering a detailed and nuanced understanding of defects. However, despite its potential for enhancing defect detection accuracy and efficiency, semantic segmentation is often constrained by the high costs associated with extensive manual annotation of data [39, 20]. A major obstacle is the requirement for extensive labeled training data which requires expert human labor.
@@
To overcome this limitation, we propose PV-S3 (Photovoltaic-Semi Supervised Segmentation), a semi-supervised deep learning framework for semantic segmentation in PV module defect detection, which efficiently utilize both labeled and unlabeled EL images. PV-S3 reduces the reliance on extensive labeled data while addressing the scalability issues in large-scale solar installations. It is based on mean-teacher approach [47] and efficiently leverages unlabeled data. The proposed method enhances model accuracy and generalization by enforcing consistency between the predictions of a student model and a temporally averaged teacher model. Such an approach is crucial for reducing reliance on extensive labeled datasets, addressing a key challenge in the semantic segmentation of PV module defects.
@@
To address the challenges posed by class imbalance and the scarcity of labeled data in real-world datasets, PV-S3 employs a novel unsupervised semi-cross entropy loss function. This loss effectively leverages both labeled and unlabeled data, enhancing the model’s ability to accurately identify rare defects. Semi-cross entropy introduces stricter positive and negative thresholds, fostering stronger convergence by deferentially weighting the contributions of various classes. This mechanism ensures that the model not only learns from the limited labeled data but also extracts valuable patterns from the unlabeled data, thereby significantly improving the detection accuracy of less frequent defect types. Figure 1 represents an high-level overview of the proposed framework for defect detection in photovoltaic cells using semi-supervised segmentation.
@@
    We study automatic defect detection in photovoltaic cells using a novel approach where we perform semantic segmentation of PV cell defects with limited annotations. This reduces the dependency on a large amount of labeled data alleviating the need for extensive manual annotation.
@@
    We propose PV-S3, a novel framework to solve this problem which can utilize both labeled and unlabeled samples to perform semantic segmentation for PV defect detection.
@@
    We develop Semi-Cross-Entropy loss, a learning objective which helps PV-S3 to detect rare defects present in limited and an imbalanced training data.
@@
    We perform extensive evaluation of our approach on several real-world PV datasets, demonstrating its effectiveness and efficiency in detecting various types of defects. The proposed approach outperforms existing state-of-the-art fully supervised methods with merely 20% annotations.
In this study, we address the challenge of employing semi-supervised learning (SSL) for automatic detection of defects in PV modules to reduce the need for extensive annotation. Our proposed method will advance the field of PV maintenance by providing an efficient, and cost-effective solution for defect detection.
@@ 2 Related Work
PV modules are essential components for converting sunlight into electrical energy, comprised of interconnected photovoltaic cells that absorb sunlight to generate direct current (DC) electricity. However, defects introduced during manufacturing, handling, and installation processes can impair their performance and reliability. Common issues include cracks, fractures, soldering problems, corrosion, and electrical contact issues [41, 45]. These defects can lead to decreased power output, hotspots, accelerated degradation, and safety risks. Timely detection and identification of these issues are crucial for ensuring optimal PV system operation and longevity. Two primary techniques used for assessing PV module performance are current-voltage (I-V) measurements and electroluminescence (EL) imaging. EL imaging is particularly valuable for maintenance evaluations and defect detection. Given the large scale of PV plants, there is a growing focus on automating defect detection and classification in EL images [49, 36, 45].
@@
Defect classification in PV modules is crucial for panel efficiency and longevity. Traditional machine learning methods like Support Vector Machines (SVMs) [42, 13], Random Forests [42, 9], and k-Nearest Neighbors (k-NN) [42, 35] have been widely applied, particularly effective for simpler defect types. However, the diverse and complex nature of PV defects demands more advanced techniques. Deep learning, especially Convolutional Neural Networks (CNNs), has gained prominence for its ability to learn from large datasets, capturing subtle defect variations [4, 13]. Modified CNN variants like multi-channel CNNs [54] and HRNet [55] have shown effectiveness in defect detection. Various CNN architectures, including VGG16 [46], ResNet50 [46], and InceptionV3 [46], have been explored for defect classification, leveraging multiple convolutional layers and transfer learning with pre-trained models for fine-tuning.
@@
Defect detection, unlike classification, demands defect localization in images, employing segmentation techniques for precise delineation. This task is vital for understanding defect characteristics and their spatial distribution in detail. Numerous models and methodologies aim for accurate segmentation of PV cell defects [20, 40, 55]. Semantic segmentation, a robust computer vision technique, has recently gained traction for this purpose, leveraging powerful pre-trained models [20, 40]. However, training segmentation models, especially for specialized domains like PV module defect detection, poses unique challenges, primarily due to extensive data annotation requirements [37]. The effectiveness of segmentation models hinges on ample labeled data availability for training, underscoring the necessity for efficient annotation strategies.
@@ 3.2.1 Supervised Loss
3.2.2 Consistency Loss
The consistency loss ensures uniformity between student and teacher network predictions by comparing their outputs on labeled and unlabeled data. In the standard mean teacher framework, Mean Square Error (MSE) typically serves as the consistency loss, aiming to minimize prediction discrepancies across datasets. However, MSE may not adequately address the complexities of defect detection in photovoltaic cells, especially concerning class imbalance and limited labeled data. To address these challenges, we introduce the SemiCE Loss tailored for semi-supervised defect detection. SemiCE enhances learning in scenarios of data scarcity and class imbalance, focusing on differential treatment of positive and negative predictions across labeled and unlabeled data.
The SemiCE loss consists of two key components: the positive loss ($\mathcal{L}_{\text{pos}}$) and the negative loss ($\mathcal{L}_{\text{neg}}$). $\mathcal{L}_{\text{pos}}$ emphasizes improving the accuracy of positive predictions crucial for defect identification. Conversely, $\mathcal{L}_{\text{neg}}$ handles negative predictions below the confidence threshold to prevent excessive noise interference. Integrating SemiCE within the mean teacher framework enhances training stability and effectiveness. By optimizing positive and negative predictions while considering unlabeled data characteristics, SemiCE fosters accurate defect detection in solar PV modules.
@@ 3.2.2 Consistency Loss
3.2.3 Augmentations
Augmentations play a crucial role in enhancing the model’s ability to detect defects in solar cells by introducing controlled variations to the training data. These variations enable the model to adapt to diverse real-world scenarios, capturing the nuances of defects under different conditions, thus contributing to improved accuracy and robustness in defect detection. PV-S3 utilizes both weak and strong augmentations to enrich the diversity of training data. Weak augmentations introduce subtle variations like random rotations and flips, enhancing the model’s capability to recognize defects across various orientations and lighting conditions. Strong augmentations, on the other hand, involve significant alterations to training images, challenging the model to adapt to complex scenarios beyond typical conditions. While weak augmentations are applied to the teacher model for generating better pseudo-labels, strong augmentations are used for training the student model, exposing it to more challenging samples to enhance its ability to generalize from complex variations, thus improving overall performance and robustness.
@@ 4 Experiments and Evaluation
We use the UCF-EL Defect Dataset, which includes nine distinct defect types, each with specific EL patterns and locations on the cell surface, such as cracks, grid interruptions, and corrosion. In Figure  3, we present a visual representation of some of the EL images containing defects. The defect are classified into four major defect classes which are Contact, Crack, Interconnect and Corrosion. Further each of these defects have sub categories for defect. Pie charts and histogram distributions of defect classes, as shown in Figures  4 and  5, indicate Contact defects are the most prevalent, followed by Crack and Interconnect, while Corrosion defects are least common.
@@
Figure 4: Distribution of defects for UCF-EL dataset: The pie chart on the left shows the distribution of 4 defect classes in the images whereas the pixel-wise pie chart on the right further details the distribution of defects at the pixel level.
@@
During the training phase, the model is trained with 20% of the labeled data and 80% of the unlabeled data from the training set. The training process involves an iterative optimization approach, where the model is trained using both labeled and unlabeled data in a semi supervised manner. The labeled data is used to compute the standard supervised loss, while the unlabeled data is employed to calculate a consistency loss, ensuring that the model produces consistent predictions for similar inputs. The training process is carried out for 30 epochs with batch size 16, Stochastic Gradient Descent (SGD) optimizer with learning rate 0.001, weight decay 0.0001 and momentum 0.9 on a 32GB GPU. Further consistency weight is taken as 1.5 and base size of image is $299\times 299$ and crop size of $224\times 224$.
@@
We evaluate the proposed method using Mean Intersection over Union (IoU), precision, recall, and F1 score. Mean IoU measures the overlap between the predicted segmentation masks and the ground truth annotations, providing an overall assessment of the segmentation accuracy. Precision represents the ratio of correctly predicted defect pixels to all predicted defect pixels, while recall measures the proportion of correctly predicted defect pixels out of all ground truth defect pixels. F1 score combines precision and recall, providing a balanced measure of the model’s overall performance.

### v3

# PV-S3111Photovoltaic-Semi Supervised Segmentation : Advancing Automatic Photovoltaic Defect Detection using Semi-Supervised Semantic Segmentation of Electroluminescence Images

@@ Abstract
Photovoltaic (PV) systems allow us to tap into all abundant solar energy, however they require regular maintenance for high efficiency and to prevent degradation. Traditional manual health check, using Electroluminescence (EL) imaging, is expensive and logistically challenging which makes automated defect detection essential. Current automation approaches require extensive manual expert labeling, which is time-consuming, expensive, and prone to errors. We propose PV-S3 (Photovoltaic-Semi Supervised Segmentation), a Semi-Supervised Learning approach for semantic segmentation of defects in EL images that reduces reliance on extensive labeling. PV-S3 is a Deep learning model trained using a few labeled images along with numerous unlabeled images. We introduce a novel Semi Cross-Entropy loss function to deal with class imbalance. We evaluate PV-S3 on multiple datasets and demonstrate its effectiveness and adaptability. With merely 20% labeled samples, we achieve an absolute improvement of 9.7% in IoU, 13.5% in Precision, 29.15% in Recall, and 20.42% in F1-Score over prior state-of-the-art supervised method (which uses 100% labeled samples) on UCF-EL dataset (largest dataset available for semantic segmentation of EL images) showing improvement in performance while reducing the annotation costs by 80%. For more details, visit our GitHub repository: https://github.com/abj247/PV-S3.
@@ 1 Introduction
The increasing global demand for clean and sustainable energy sources has highlighted the importance of solar energy, particularly photovoltaic (PV) systems, in the renewable energy landscape [20, 23, 38]. While PV modules are essential for converting sunlight into electricity, their efficiency and reliability are often compromised by various defects, ranging from manufacturing imperfections to environmental stressors [19, 28, 16, 26]. These defects reduce the overall efficiency and the life cycle while impacting the financial viability of PV systems. Various imaging techniques such as Infrared Thermography (IRT), Photoluminescence (PL) Imaging, and Electroluminescence (EL) imaging are currently used, however, EL imaging remains a key technique due to its high resolution and detailed electrical property information [43, 25, 4, 10, 45]. Traditionally, defect detection in PV modules has relied on manual inspection of EL images [5, 14, 13]. Manual inspection however requires certain expertise while being labor-intensive and prone to human error. Recent expansion of PV installations increased the number of images that need to be analyzed thus exacerbating the problem. In long run automation is the key to address all these issues.
@@
Application of artificial intelligence (AI) to bring automation has been widely accepted in the energy industry [9, 22] and owing to above challenges, application for automated defect detection in PV modules has also gained attention [1, 2, 3, 40]. AI-based approaches, including binary classification for defective/non-defective, multi-class classification for identifying types of defects, and object detection for locating defects, pave the way for comprehensive analysis of PV systems [21, 52]. Among these, semantic segmentation stands out for its ability to delineate and classify each pixel of an image according to defect type, offering a detailed and nuanced understanding of defects. However, despite its potential for enhancing defect detection accuracy and efficiency, semantic segmentation is often constrained by the high costs associated with extensive manual annotation of data [33, 17]. A major obstacle is the requirement for extensive labeled training data which requires expert human labor.
@@
To overcome this limitation, we propose PV-S3 (Photovoltaic-Semi Supervised Segmentation), a semi-supervised deep learning framework for semantic segmentation in PV module defect detection, which efficiently utilize both labeled and unlabeled EL images. PV-S3 reduces the reliance on extensive labeled data while addressing the scalability issues in large-scale solar installations. It is based on mean-teacher approach [42] and efficiently leverages unlabeled data. The proposed method enhances model accuracy and generalization by enforcing consistency between the predictions of a student model and a temporally averaged teacher model. Such an approach is crucial for reducing reliance on extensive labeled datasets, addressing a key challenge in the semantic segmentation of PV module defects. To the best of our knowledge there is no current existing work in Photovoltaics making use of semi-supervised learning approaches or reduced labeled data for defect detection.
@@
To address the challenges posed by the scarcity of labeled data in real-world datasets, we propose a semi-supervised approach. We effectively leverages both labeled and unlabeled data to enhance the model’s ability to detect defects. Our Semi-cross entropy loss introduces stricter positive and negative thresholds, fostering stronger convergence by deferentially weighting the contributions of various classes. This ensures that the model extracts valuable patterns from the unlabeled data. Figure 1 represents an high-level overview of the proposed framework for defect detection in photovoltaic cells using semi-supervised segmentation.
@@
    We study the novel problem of automatic defect detection in photovoltaic (PV) modules under limited annotations, reducing the dependency on large labeled datasets and alleviating the need for extensive manual annotation—a problem not previously explored in existing literature.
@@
    We propose PV-S3, a semi-supervised framework for semantic segmentation of PV modules, which effectively leverages both labeled and unlabeled data to enhance defect detection accuracy while maintaining label efficiency.
@@
    We perform extensive evaluation of the proposed approach on several real-world PV datasets, demonstrating its effectiveness and efficiency in detecting various types of defects.
@@
    We demoonstrate that the proposed approach outperforms existing state-of-the-art fully supervised methods with merely 20% annotations.
In this study, we employ semi-supervised learning (SSL) for automatic detection of defects in PV modules to reduce the need for extensive annotation. The proposed method will advance the field of PV maintenance by providing an efficient, and cost-effective solution for defect detection.
@@ 2 Related Work
PV modules are essential components for converting sunlight into electrical energy, comprised of interconnected photovoltaic cells that absorb sunlight to generate direct current (DC) electricity. However, defects introduced during manufacturing, handling, and installation processes can impair their performance and reliability. Common issues include cracks, fractures, soldering problems, corrosion, and electrical contact issues [35, 41]. These defects can lead to decreased power output, hotspots, accelerated degradation, and safety risks. Timely detection and identification of these issues are crucial for ensuring optimal PV system operation and longevity. Two primary techniques used for assessing PV module performance are current-voltage (I-V) measurements and electroluminescence (EL) imaging. EL imaging is particularly valuable for maintenance evaluations and defect detection. Given the large scale of PV plants, there is a growing focus on automating defect detection and classification in EL images [44, 31].
@@
Defect classification in PV modules is crucial for panel efficiency and longevity. Traditional machine learning methods like Support Vector Machines (SVMs) [37], Random Forests [37, 8], and k-Nearest Neighbors (k-NN) [37] along with ensemble learning techniques [36] have been widely applied, particularly effective for simpler defect types. However, the diverse and complex nature of PV defects demands more advanced techniques. Deep learning, especially Convolutional Neural Networks (CNNs), has gained prominence for its ability to learn from large datasets, capturing subtle defect variations [30]. Modified CNN variants like multi scale CNNs with transfer learning [27] and multi-scale residual CNNs with class balancing strategy [47] have shown effectiveness in defect detection. Recently vision transformer based architectures have also been used for defect detection if PV Cells [15]. YOLO based architecture have also been used to detect defects in the electroluminescence images for PV defect detection. YOLOv8 improved architecture is used to get enhanced performance [7]. Adaptive fusion integrated with YOLO architectures to expand the backbone feature’s receptive field and adaptively filter out conflicting information across varied hierarchical levels improved the detection performance [50].
@@
Defect detection, unlike classification, demands defect localization in images, employing segmentation techniques for precise delineation. This task is vital for understanding defect characteristics and their spatial distribution in detail. Numerous models and methodologies aim for accurate segmentation of PV cell defects [17, 34]. Semantic segmentation, a robust computer vision technique, has recently gained traction for this purpose, leveraging powerful pre-trained models [17, 34]. However, training segmentation models, especially for specialized domains like PV module defect detection, poses unique challenges, primarily due to extensive data annotation requirements [32]. The effectiveness of segmentation models hinges on ample labeled data availability for training, underscoring the necessity for efficient annotation strategies.
@@ 3.2.1 Supervised Loss
3.2.2 SemiCE as Consistency Loss
The consistency loss ensures uniformity between student and teacher network predictions by comparing their outputs on labeled and unlabeled data. We use SemiCE Loss as consistency loss adapted for semi-supervised defect detection. SemiCE enhances learning in scenarios of data scarcity and class imbalance, focusing on differential treatment of positive and negative predictions across labeled and unlabeled data. The SemiCE loss consists of two key components: the positive loss ($\mathcal{L}_{\text{pos}}$) and the negative loss ($\mathcal{L}_{\text{neg}}$). $\mathcal{L}_{\text{pos}}$ emphasizes improving the accuracy of positive predictions crucial for defect identification. Conversely, $\mathcal{L}_{\text{neg}}$ handles negative predictions below the confidence threshold to prevent excessive noise interference. SemiCE loss enhances training stability and effectiveness by specifically addressing the challenges posed by class imbalance. By optimizing positive and negative predictions while considering unlabeled data characteristics, SemiCE aims on accurate defect detection in PV modules.
@@ 3.2.2 SemiCE as Consistency Loss
3.3 Data Pre-Processing and Augmentations
Data Pre-Processing and augmentations are crucial in enhancing our model’s ability to detect defects by introducing controlled variations into the training data. These variations enable the model to adapt to diverse real-world scenarios, capturing the nuances of defects under different conditions, thus contributing to improved accuracy and robustness in defect detection. Image pre-processing involves resizing and scaling pixel values as per the dataset and model requirements. They include normalizing images and converting images to tensors suitable for neural network input. In our approach, we utilize both weak and strong augmentations to enrich the diversity of the training data.Weak augmentations are applied to the teacher model to generate better pseudo-labels. These subtle changes ensure the teacher model produces reliable pseudo-labels while being exposed to slight variations in the data. Strong augmentations are used for training the student model, exposing it to more challenging samples to enhance its ability to generalize from complex variations. They involve significant alterations to training images such as applying color jittering, random grayscale conversion and blurring the images, and also cut mix augmentation [51]. These augmentation techniques improve the overall performance and robustness of the model. Our semi-supervised learning framework and augmentation can be adapted to other datasets with the same pre-processing steps performed on the studied datasets. In this study, we have adapted the proposed model to two different datasets. This adaptability enhances the reproducibility and generalizability of our approach across various datasets.
@@ 4 Experiments and Evaluation
We use the UCF-EL Defect Dataset, which includes nine distinct defect types, each with specific EL patterns and locations on the cell surface, such as cracks, grid interruptions, and corrosion. The dataset also consists of images which are healthy and contain no defects. The ground truth segmentation map for these images has a pixel index same as the background.
In Figure  3, we present a visual representation of some of the EL images containing defects. The defect are classified into four major defect classes which are Contact, Crack, Interconnect and Corrosion. Further each of these defects have sub categories for defect. Pie charts and histogram distributions of defect classes, as shown in Figures  4 and  5, indicate Contact defects are the most prevalent, followed by Crack and Interconnect, while Corrosion defects are least common.
@@
Figure 4: Distribution of defects for the UCF-EL dataset: The pie chart on the left illustrates the distribution of the four defect classes based on the number of images containing each defect class, indicating how frequently each type of defect appears in the dataset at the image level. In contrast, the pie chart on the right shows the pixel-wise distribution of defects, detailing the proportion of pixels occupied by each defect class across all images. This comparison highlights the difference between the occurrence of defects in images and their actual spatial extent within the dataset.
@@
During the training phase, the model is trained with 20% of the labeled data and 80% of the unlabeled data from the training set. The training process is carried out for 30 epochs with batch size 16, Stochastic Gradient Descent (SGD) optimizer with learning rate 0.001, weight decay 0.0001 and momentum 0.9 on a 32GB GPU. Further consistency weight is taken as 1.5 and base size of image is $299\times 299$ and crop size of $224\times 224$. Further strong and weak augmentations are applied to student and teacher model. Strong augmentations include normalizing imags, color jittering (adjusting brightness, contrast, saturation, and hue by up to ±$\pm 50\%$), random grayscale conversion with a 20% probability, Gaussian blur with a random kernel size between 1 and 5 and sigma between 0.1 and 2.0, and CutMix augmentation where patches of images are randomly replaced with patches from other images based on a beta distribution ($\alpha=4$, $\beta=4$). Weak augmentations applied to teacher network only include normalizing the images. The standard mean and standard deviation values of $[0.485,0.456,0.406]$ and $[0.229,0.224,0.225]$ respectively are used for normalizing the images for both student and teacher models.
@@
We evaluate the proposed method using Mean Intersection over Union (IoU), precision, recall, and F1 score. Mean IoU measures the overlap between the predicted segmentation masks and the ground truth annotations, providing an overall assessment of the segmentation accuracy. Precision represents the ratio of correctly predicted defect pixels to all predicted defect pixels, while recall measures the proportion of correctly predicted defect pixels out of all ground truth defect pixels. F1 score combines precision and recall, providing a balanced measure of the model’s overall performance. To clearly define these evaluation metrics, we provide their mathematical formulations:
  • 1.
    Precision is calculated as:
    Precision=True Positives (TP)True Positives (TP)+False Positives ($\text{Precision}=\frac{\text{True Positives (TP)}}{\text{True Positives (TP)}+\text{False Positives (FP)}}$ (6)
    where TP represents the number of correctly predicted positive pixels, and FP represents the number of pixels incorrectly predicted as positive.
  • 2.
    Recall is calculated as:
    Recall=True Positives (TP)True Positives (TP)+False Negatives ($\text{Recall}=\frac{\text{True Positives (TP)}}{\text{True Positives (TP)}+\text{False Negatives (FN)}}$ (7)
    where FN represents the number of positive pixels that were not correctly predicted.
  • 3.
    F1 Score is the harmonic mean of precision and recall:
[…]
  • 4.
    Mean Intersection over Union (IoU) is calculated as:
    Mean $\text{Mean IoU}=\frac{1}{N_{\text{classes}}}\sum_{i=1}^{N_{\text{classes}}}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}$ (9)
    where $N_{\text{classes}}$ is the number of classes, and $\text{TP}_{i}$, $\text{FP}_{i}$, and $\text{FN}_{i}$ are the true positives, false positives, and false negatives for class $i$, respectively.
Here, true positives (TP) are pixels correctly identified as defects, false positives (FP) are pixels incorrectly identified as defects, and false negatives (FN) are defect pixels that were missed by the model. These equations provide a quantitative basis for evaluating and comparing the performance of our segmentation model.